### v2

# CuClarabel: GPU Acceleration for a Conic Optimization Solver

Abstract
We present the GPU implementation of the general-purpose interior-point solver Clarabel for convex optimization problems with conic constraints. We introduce a mixed parallel computing strategy that processes linear constraints first, then handles other conic constraints in parallel. This mixed parallel computing strategy currently supports linear, second-order cone, exponential cone, and power cone constraints. We demonstrate that integrating a mixed parallel computing strategy with GPU-based direct linear system solvers enhances the performance of GPU-based conic solvers, surpassing their CPU-based counterparts across a wide range of conic optimization problems. We also show that employing mixed-precision linear system solvers can potentially achieve additional acceleration without compromising solution accuracy.
1 Introduction
We consider the following convex optimization problem [1] with a quadratic objective and conic constraints:
$\begin{array}[]{ll}\text{minimize}&\frac{1}{2}x^{T}Px+q^{T}x\\ \text{subject to}&Ax+s=b,\\ &s\in\mathcal{K},\end{array}$ ($\mathcal{P}$)
[…]
The dual problem of ($\mathcal{P}$) is
$\begin{array}[]{ll}\text{maximize}&-\frac{1}{2}x^{T}Px-b^{T}z\\ \text{subject to}&Px+A^{T}z=-q,\\ &z\in\mathcal{K}^{*},\end{array}$ ($\mathcal{D}$)
with respect to $x,z$ and where $\mathcal{K}^{*}$ is the dual cone of $\mathcal{K}$. Solving $\mathcal{P}$ and $\mathcal{D}$ is equivalent to solving the Karush-Kuhn-Tucker (KKT) conditions when strong duality holds. On the other hand, the set of strongly primal infeasibility certificates for ($\mathcal{P}$) is
$\mathbb{P}$ =$=\left\{z\ \left|\ A^{T}z=0,~{}z\in\mathcal{K}^{*},~{}\left\langle{b},{z}\right\rangle<0\right.\right\},$ (1)
and the set of strongly dual infeasibility certificates is
$\mathbb{D}$ =$=\left\{x\ \left|\ ~{}Px=0,~{}-Ax\in\mathcal{K},~{}\left\langle{q},{x}\right\rangle<0\right.\right\}.$ (2)
Finding an optimal solution with strong duality or detecting a strongly infeasible certificate can be unified into solving a linear complementarity problem with two additional slack variables $\tau,\kappa\geq 0$ [19], which can be reformulated to the following problem [15]:
$\begin{array}[]{ll}\text{minimize}&s^{T}z+\tau\kappa\\ \text{subject to}&\frac{1}{\tau}x^{T}Px+q^{T}x+b^{T}z=-\kappa,\\ &Px+A^{T}z+q\tau=0,\\ &Ax+s-b\tau=0,\\ &(s,z,\tau,\kappa)\in\mathcal{K}\times\mathcal{K^{*}}\times\textbf{R}_{+}\times\textbf{R}_{+},\end{array}$ ($\mathcal{H}$)
[…]
  1. i)
    If $\tau^{\star}>0$ then $(x^{\star}/\tau^{\star},s^{\star}/\tau^{\star})$ is an optimal solution to ($\mathcal{P}$) and $(x^{\star}/\tau^{\star},z^{\star}/\tau^{\star})$ is an optimal solution to $(\mathcal{D})$.
  2. ii)
    If $\kappa^{\star}>0$ then at least one of the following holds:
    • ($\mathcal{P}$) is strongly infeasible and $z^{\star}\in\mathbb{P}$.
    • $(\mathcal{D})$ is strongly infeasible and $x^{\star}\in\mathbb{D}$.
The optimal solution $\tau^{\star},\kappa^{\star}$ satisfy the complementarity slackness condition, i.e., at most one of $\tau^{\star},\kappa^{\star}$ is nonzero. The pathological case $\tau^{\star}=\kappa^{\star}=0$ has been discussed in [19].
[…]
1.1 Our contribution
We describe a Julia GPU implementation of the general-purpose interior-point solver Clarabel [15]. In our implementation, we support cases where $\mathcal{K}$ is an intersection of atomic cones: zero cones, nonnegative cones, second order cones, exponential cones, and power cones. We propose a mixed parallel computing strategy that parallelizes computing for each type of cone, integrates the CUDSS [21] library for linear system solving, and supports mixed-precision linear system solves for moderate speed improvements. Furthermore, we evaluate our solver against others across a variety of conic optimization problems. Our implementation of CuClarabel is available at https://github.com/cvxgrp/CuClarabel.
1.2 Related work
Interior-point methods and solver development.
The interior-point method [20] was first discovered by Dikin [22], and became more mainstream after the conception of Karmarkar’s method [23], a polynomial-time algorithm for linear programming, and Renegar’s [24] path following method. Interior-point methods are known for their ability to solve conic optimization problems to high precision, and is chosen as the default algorithm for many conic optimization solvers [12, 25, 15, 26, 27]. Common variations of the interior-point method include potential reduction methods and path-following methods.
Interior-point methods employ a Newton-like strategy to compute a search direction at every iteration. Unfortunately, the matrix factorization required to compute this direction which scales with the dimension of an optimization problem, rendering very large problems difficult to solve. Developing efficient interior-point methods on exotic cones directly is a promising research direction [26, 27] to alleviate this computational burden. Using exotic cones, we can represent equivalent problems with significantly fewer variables and exploit sparse structure within these exotic cones for efficient implementation of interior-point methods [25, 28, 29]. However, operations within an exotic cone can hardly be parallelized, while parallelism across heterogeneous cones of different dimensionalities will introduce significant synchronization delay.
GPU acceleration in optimization algorithms.
GPUs are playing an increasingly significant role in scientific computing. In optimization, GPUs are used to run solvers based on first-order methods. For example, CuPDLP [30] is based on the popular PDHG algorithm [31], which requires only matrix multiplication and addition without the use of direct methods (i.e. it is factorization-free). Solvers that require the solution to linear systems, like SCS [13] and CuOSQP [32], have relied on indirect iterative methods—such as the conjugate gradient (CG), the minimal residual (MINRES) [33], and the generalized minimum residual (GMRES) [34] methods—to solve these systems on GPUs. However, the linear systems solved in first-order methods are generally much better conditioned than those encountered in interior-point methods, where the linear systems become increasingly ill-conditioned as the iterations progress. As an interior-point method approaches higher precision, the number of iterations for each inner indirect linear solves increases significantly, which will eventually offset benefits of GPU parallelism and make GPU-based interior-point methods less preferable compared to CPU-based solvers with direct methods in overall computational time [35]. Recently, NVIDIA released the CUDSS package [36], which provides fast direct methods on GPUs for sparse linear systems. Previous work has integrated CUDSS in a nonlinear optimization solver [37], resulting in significant speed-up on large-scale problems compared to its CPU-based counterpart.
Mixed-precision methods.
Mixed-precision, or multiprecision, methods [38, 39] have become progressively more popular due to their synergies with modern GPU architectures. For example, in lower-precision configuration one enjoys significant speedup in algorithms, as modern architectures have 32-bit implementations that are around twice as fast as their 64-bit counterparts [38, 40]. Mixed precision methods aim to capitalize on the computational benefits of performing expensive operations in lower precision, while maintaining (or reducing the negative impact to) the superior numerical accuracy from higher precision. Classically, mixed-precision methods for direct (linear system solving) methods involve factorizing a matrix in lower precision, and then applying iterative refinement. In this approach, only the more expensive steps, such as the matrix factorizations and backsolves, are done in lower precision. Mixed-precision methods are used throughout scientific computing, with applications including BLAS operations [41], different linear system solvers such as Krylov methods (CG, GMRES) [39, 42, 43], solving partial differential equations [44], and training deep neural networks [44, 45, 46].
1.3 Paper outline
In §2, we review and discuss the interior point method used by Clarabel [15] for solving conic optimization problems. In §3, we outline how to implement parallel computation for cone operations and solve linear systems within our GPU solver. §4 details our numerical experiments. We detail the supported cones in §A, how to calculate the barrier functions in §B and scaling matrices for each cone in §C.
2 Overview of Clarabel
We first sketch briefly the main operations used in the Clarabel solver [15] for computing a solution to ($\mathcal{P}$). Solving the equivalent problem ($\mathcal{H}$) amounts to finding a root of the following nonlinear equations
$G(x,z,s,\tau,\kappa)=\begin{bmatrix}0\\ s\\ \kappa\end{bmatrix}-\begin{bmatrix}\hphantom{+}P&A^{T}&q\\ -A&0&b\\ -q^{T}&-b^{T}&0\end{bmatrix}\begin{bmatrix}x\\ z\\ \tau\end{bmatrix}+\begin{bmatrix}0\\ 0\\ \frac{1}{\tau}x^{T}Px\end{bmatrix}=0,$ (3)
$(x,z,s,\tau,\kappa)\in\mathcal{F}=\textbf{R}^{n}\times\mathcal{K}\times\mathcal{K}^{*}\times\textbf{R}_{+}\times\textbf{R}_{+}.$
Besides the zero cone that is a linear constraint, other supported conic constraints are smoothed by nonlinear equations in pairs within an interior-point method,
$s=-\mu\nabla f(z),\quad\tau\kappa=\mu,$ (4)
where $\mu$ is the smoothing parameter and $f(\cdot)$ is the logarithmically homogeneous self-concordant barrier (LHSCB) function for cone $\mathcal{K}^{*}$. The barrier functions for different cones are detailed in §B. The trajectory (also called the central path)
[…]
where $v=(x,z,s,\tau,\kappa)$, characterizes the solution of (3) in the right limit $\mu\rightarrow 0$, given an initial point $v^{0}$. After starting from $v^{0}$, the Clarabel solver iterates the following steps for each iteration $k$.
Update residuals and check the termination condition.
We update several key metrics at the start of each iteration. Defining the normalized variables $\bar{x}=x/\tau,\bar{s}=s/\tau,\bar{z}=z/\tau$, the primal and dual residuals are then
$r_{p}$ $=A\bar{x}-\bar{s}+b,$
$r_{d}$ $=P\bar{x}+A^{\top}\bar{z}+q,$
with corresponding primal and dual objectives
$g_{p}$ $=\frac{1}{2}\bar{x}^{\top}P\bar{x}+q^{T}x,$
$g_{d}$ $=-\frac{1}{2}\bar{x}^{\top}P\bar{x}-b^{T}\bar{z},$
and complementarity slackness $\mu=\frac{{s}^{\top}z+\kappa\tau}{\nu+1}$. The solver returns an approximate optimal point if
$\left\|r_{p}\right\|_{\infty}$ $<\epsilon_{\text{feas}}\max\{1,\left\|b\right\|_{\infty}+\left\|\bar{x}\right\|_{\infty}+\left\|\bar{s}\right\|_{\infty}\}$
$\left\|r_{d}\right\|_{\infty}$ $<\epsilon_{\text{feas}}\max\{1,\left\|q\right\|_{\infty}+\left\|\bar{x}\right\|_{\infty}+\left\|\bar{z}\right\|_{\infty}\}$
$|g_{p}-g_{d}|$ $<\epsilon_{\text{feas}}\max\{1,\min\{|g_{p}|,|g_{d}|\}\}.$
Otherwise, it returns a certificate of primal infeasibility if
$\left\|A^{T}z\right\|_{\infty}$ $<-\epsilon_{\text{inf}}\max(1,\left\|x\right\|_{\infty}+\left\|z\right\|_{\infty})(b^{T}z)$
$b^{T}z$ $<-\epsilon_{\text{inf}},$
and a certificate of dual infeasibility if
$\left\|Px\right\|_{\infty}$ $<-\epsilon_{\text{inf}}\max(1,\left\|x\right\|_{\infty})(b^{T}z)$
$\left\|Ax+s\right\|_{\infty}$ $<-\epsilon_{\text{inf}}\max(1,\left\|x\right\|_{\infty}+\left\|s\right\|_{\infty})(q^{T}x)$
$q^{T}x$ $<-\epsilon_{\text{inf}}.$
Note that $\epsilon_{\text{feas}},\epsilon_{\text{inf}}$ are predefined parameters within the Clarabel solver.
Find search directions.
We then compute Newton-like search directions using a linearization of the central path (5). In other words, we solve the following linear system given some right-hand side residual $d=(d_{x},d_{z},d_{\tau},$ $d_{s},d_{\kappa})$,
$\begin{bmatrix}0\\ \Delta s\\ \Delta\kappa\end{bmatrix}-\begin{bmatrix}P&A^{T}&q\\ -A&0&b\\ -(q+2P\xi)^{T}&-b^{T}&\xi^{T}P\xi\end{bmatrix}\begin{bmatrix}\Delta x\\ \Delta z\\ \Delta\tau\end{bmatrix}=-\begin{bmatrix}d_{x}\\ d_{z}\\ d_{\tau}\end{bmatrix}$ (6a)
$H\Delta z+\Delta s=-d_{s},\quad\kappa\Delta\tau+\tau\Delta\kappa=-d_{\kappa},$ (6b)
where $H$ is the same scaling matrix as in the CPU version of Clarabel [15] and described in §C. We have shown in [15] that solving (6) reduces to solve the next linear system with two different right-hand sides,
$\underbrace{\begin{bmatrix}P&A^{\top}\\ A&-H\end{bmatrix}}_{K}\left[\begin{array}[]{c|c}\Delta x_{1}&\Delta x_{2}\\ \Delta z_{1}&\Delta z_{2}\end{array}\right]=\left[\begin{array}[]{c|c}d_{x}&-q\\ -(d_{z}-d_{s})&b\end{array}\right].$ (7)
After solving (7), we recover the search direction $\Delta=(\Delta x,\Delta z,\Delta\tau,\Delta s,\Delta\kappa)$ using
$\Delta\tau=\frac{d_{\tau}-d_{\kappa}/\tau+(2P\xi+q)^{\top}\Delta x_{1}+b^{\top}\Delta z_{1}}{\kappa/\tau+\xi^{\top}P\xi-(2P\xi+q)^{\top}\Delta x_{2}-b^{\top}\Delta z_{2}}$
$=\frac{d_{\tau}-d_{\kappa}/\tau+q^{\top}\Delta x_{1}+b^{\top}\Delta z_{1}+2\xi^{\top}P\Delta x_{1}}{\|\Delta x_{2}-\xi\|_{P}^{2}-\|\Delta x_{2}\|_{P}^{2}-q^{\top}\Delta x_{2}-b^{\top}\Delta z_{2}},$ (8a)
and
$\Delta x=\Delta x_{1}+\Delta\tau\Delta x_{2},\quad\Delta z=\Delta z_{1}+\Delta\tau\Delta z_{2},$ (8b)
$\Delta s=-d_{s}-H\Delta z,\quad\Delta\kappa=-(d_{\kappa}+\kappa\Delta\tau)/\tau.$ (8c)
In an interior-point method with a predictor-corrector scheme, we need to solve (6) with two different values for $d$. The first is for the affine step (predictor) with
$d=(G(x,z,s,\tau,\kappa),\kappa\tau,s).$
The other assigns $d$ as
$(d_{x},d_{z},d_{\tau})=(1-\sigma)G(x,z,s,\tau,\kappa),\quad d_{\kappa}=\kappa\tau+\Delta\kappa\Delta\tau-\sigma\mu,\\ d_{s}=\left\{\begin{array}[]{cl}W^{\top}\left(\lambda\backslash\left(\lambda\circ\lambda+\eta-\sigma\mu\mathbf{e}\right)\right)&\text{(symmetric)}\\ s+\sigma\mu\nabla f(z)+\eta&\text{(nonsymmetric),}\end{array}\right.\end{aligned}$ (9)
[…]
$\eta=(W^{-1}\Delta s)\circ(W\Delta z),$ (10a)
for symmetric cones and the $3$rd-order correction [50]
$\eta=-\frac{1}{2}\nabla^{3}f(z)[\Delta z^{a},\nabla^{2}f(z)^{-1}\Delta s^{a}],$ (10b)
for nonsymmetric cones. The centering parameter $\sigma$ controls the decreasing speed of both the residual $G(x,z,s,\tau,\kappa)$ and the complementarity slackness $\mu$. which is determined heuristically by value of the affine step size $\alpha_{a}$, which is the maximal value ensuring $v+\alpha_{a}\Delta_{a}\in\mathcal{F}$.
Update iterates.
At the end of each iteration $k$, we move the current iterate $v$ along the combined direction $\Delta_{c}$ and obtain the new iterate $v+\alpha_{c}\Delta_{c}$. The combined step size $\alpha_{c}$ should satisfy $v+\alpha_{c}\Delta_{c}\in\mathcal{F}$, and furthermore we have to ensure that the new iterate $v+\alpha_{c}\Delta_{c}$ stays in the neighborhood of the central path (5[50].
3 GPU formulation of Clarabel
3.1 A primer on GPU programming
Originally developed for computationally demanding gaming applications, graphics processing units (GPUs) are massively parallel, multithreaded, manycore processors, with massive computational power and memory bandwidth. As a result, GPUs are used throughout scientific computing. In this section, we outline some core properties of GPU computing devices to highlight why a GPU implementation is a natural extension to Clarabel. We then discuss our specific implementation details.
Due to their natural parallelism, GPUs differ dramatically from CPUs in the way their transistors are configured. GPUs have a smaller number of caches (blocks where memory access is fast) and instruction processing blocks, and far more, albeit simpler, computational blocks (i.e. arithmetic logic and floating point units). This suggests that CPUs utilize their larger caches to minimize instruction and memory latency within each thread, while GPUs switch between their significantly larger number of threads to hide said latency.
GPUs use a single instruction, many threads (SIMT) approach. In practice, the GPU receives a stream of instructions – each instruction is sent to groups of GPU cores (also known as a warp) and acts on multiple data in parallel. Each group of GPU cores, as a result, has single instruction, many data (SIMD) structure. This contrasts the traditional CPU vector lane approach of single instruction, single data. Note that modern CPUs also support SIMD, but at a much smaller scale than GPUs, as implied by the transistor layout. Furthermore, the SIMD structure requires data parallelism for parallel execution on a GPU.
In our approach, we utilize the SIMD structure of the GPU to accelerate our interior-point method. For each iteration of our solving algorithm, we execute the following steps:
  Interior-point method. Main computing steps.
  1.    1.
    Update the variable $v$.
  2.    2.
    Update residuals and objective values for termination check.
  3.    3.
    Update the scaling matrix $H$.
  4.    4.
    Factorize the matrix $K$ in (7).
  5.    5.
    Compute the right-hand sides of (7) for the constant, affine step, and combined steps.
  6.    6.
    Solve the linear systems (7) three times with different right-hand sides.
  7.    7.
    Recover the step direction $\Delta$ via (8).
  8.    8.
    Perform a line search and neighborhood check with respect to the central path.
The matrix factorization (step 4) and the back-solve (step 6) are the most time-consuming parts in each iteration of an interior point method [48] and can be computed using the CUDSS package. Steps 1 and 2 contain matrix addition and multiplication operations that have already been supported in CUDA, while steps 3, 5, 7, and 8 include cone operations that should be tailored for each cone. We will detail how to parallelize each of these steps in the next subsection.
3.2 Mixed parallel computing strategy
The cone operations related to steps 3, 5, 7, and 8 above require only information local to each constituent cone, and hence can be executed concurrently with respect to individual cones. We propose the mixed parallel computing strategy for a cone operation across different types of cones. Each family of cones is handled in parallel, and families of cones are handled sequentially.
As stated earlier, the cone operations related to zero cones and nonnegative cones are simply vector additions and multiplications, which have already been parallelized in CUDA. In our implementation we aggregate all zero cones into a single zero cone in our preprocessing step. The same holds for nonnegative cones. For the remaining cones, we parallelize within each family of cones (i.e., second-order cone, exponential cone, etc.). For each family of cones, we allocate a thread to each cone belonging to that family, and execute in parallel. We synchronize threads after executing for each type of cone. This adheres to the SIMD computing paradigm, as each type of cone has its own set of instructions for updating its scaling matrix.
Note that the SIMD GPU computing structure naturally favors balanced workloads, i.e., the workloads in each thread should be similar so that the synchronization will not take too much time. The exponential and power cones are 3-dimensional nonsymmetric cones that all cone operations have balanced workload among the same class of cones, thus we can parallelize the computation for each cone. On the other hand, the second-order cones may vary in dimensionality, which may not mesh well with the SIMD computing paradigm. In most use cases for large-scale second-order cone problems, e.g. optimal power flow problems [2] and finite-element problems [5], the second-order cones are of small (less than 5) dimensionality; thus the effect of this workload imbalance is negligible. Regardless, we support second-order cones of all sizes. In the next section, we outline a preprocessing procedure for our second-order cones we use to balance workloads.
We illustrate our mixed parallel computing strategy via step 3, the scaling matrix update. Recall that the conic constraint $\mathcal{K}$ can be decomposed as a Cartesian product of $p$ constituent atomic cones $\mathcal{K}_{1}\times\dots\times\mathcal{K}_{p}$ that are ordered by cone types. We assume $\mathcal{K}_{1}$ is a zero cone, $\mathcal{K}_{2}$ is a nonnegative cone, $\mathcal{K}_{3}$ to $\mathcal{K}_{i}$ are second-order cones, $\mathcal{K}_{i+1}$ to $\mathcal{K}_{j}$ are exponential cones and $\mathcal{K}_{j+1}$ to $\mathcal{K}_{p}$ are power cones. Due to our composition into constituent atomic cones, $H$ is a block-diagonal matrix,
$H=\begin{bmatrix}H_{1}&&\\ &\ddots&\\ &&H_{p}\end{bmatrix},$
where $H_{1}$ and $H_{2}$ are diagonal matrices and each block $H_{t},t\geq 3$ is the scaling matrix corresponding to cone $\mathcal{K}_{t}$ of small dimensionality.
For the update of $H$ at iteration $k+1$, we first set diagonal terms of the scaling matrix $H_{1}^{k+1}$ for the zero cone to all 00s, and then update the diagonals of the scaling matrix $H_{2}^{k+1}$ for the nonnegative cone by element-wise vector division, as in the Nesterov-Todd (NT) scaling [51]. These computations are already parallelized by CUDA. For conic constraints other than the zero cone and the nonnegative cone, the corresponding parts in the scaling matrix $H$ are no longer diagonal, but we observe $H_{t}$ only requires local information within each constituent cone, and thus we can solve for each $H_{t}$ concurrently, independent to the other cones. We implement kernel functions _soc_update_H(), _exp_update_H() and _pow_update_H() for second-order cones, exponential cones and power cones respectively. The kernel function will process one cone operation per thread, and the kernel functions belongs to the same class of cones are executed in parallel and finally synchronized by calling @sync. Details regarding the update functions for each type of cone and Algorithm 1 illustrating the update of scaling matrix $H$ can also be found in §C.
Since the mixed parallel computing strategy is independent of choices of optimization algorithms, it is also applicable for GPU implementations for cone operations in first-order operator-splitting conic solvers, like SCS [13] and COSMO [14].
3.3 Second-order cone preprocessing for balanced workloads
We can decompose any large second-order cone into a set of small second-order cones of fixed dimensionality $d$. The workload is then balanced for each second-order cone, adhering to the SIMD computing paradigm. For an $n$-dimensional second-order cone $\mathcal{K}_{soc}^{n}\{x\in\textbf{R}^{n}\mid x_{1}\geq\sqrt{\sum_{i=2}^{n}x_{i}^{2}}\}$, we can decompose it into the equivalent set of $q=\lceil(n-2)/(d-2)\rceil$ second-order cones:
$\begin{array}[]{cl}x_{1}&\geq\sqrt{x_{2}^{2}+\cdots+x_{d-1}^{2}+z_{1}^{2}}\\ z_{1}&\geq\sqrt{x_{d}^{2}+\cdots+x_{2d-2}^{2}+z_{2}^{2}}\\ &\vdots\\ z_{q}&\geq\sqrt{x_{n-d}^{2}+\cdots+x_{n}^{2}}\end{array},$
where $z\in\textbf{R}^{q}$ is an a slack variable. Each of the first $q-1$ second-order cones are $d$-dimensional, while the $g^{\textrm{th}}$ cone has dimension not larger than $d$. The workload is now balanced among these $p$ cones.
3.4 Data structures
Since data parallelism is required for parallel execution on GPUs, we manage data for each cone as a structure of arrays (SoA) in our GPU implementation, in contrast to the existing CPU counterpart which uses an array of structures (AoS). In other words, instead of creating structs for $\mathcal{K}_{i},i=1,\dots,p$ and storing pointers to each struct in an array, we concatenate the same type of local variable from different cones into a global variable and then store it in a global struct for $\mathcal{K}$, which is illustrated in Figure 1.
[…]
Figure 1: Illustration of AoS (CPU) and SoA (GPU) data structure
In addition, indexing cones in order in the setup phase can simplify the implementation of mixed parallel computing strategy. We reorder the input cones such that memory is coalescing for the same class of cones, which can accelerate computation on GPUs. Matrices are stored in the compressed sparse row (CSR) format. Instead of storing only the triangular part of a square matrix as in the CPU-based Clarabel, we store the full matrix for more efficient multiplication on the GPU.
3.5 Solving linear systems
Most of the computation time for our interior point method is spent in factorizing the matrix $K$ and in the three backsolve operations in (7). Our GPU implementation also leverages the power of the newly released sparse linear system solver CUDSS [36] for the $LDL^{T}$ factorization and backsolve operations. The iterative refinement [52] is implemented to increase the numerical stability of the backsolve operation.
Currently, a GPU has more computing cores for Float32 than Float64 and hence better performance for parallel algorithms. Also, Float32 requires less memory and takes less time for the same computation than Float64. However, lowering the precision will introduce numerical instability for solving linear systems. Thus, we employ a mixed precision solve in our matrix solves. Specifically, we use mixed precision for data within the iterative refinement step:
  1. 1.
    Solve for the residual at step $i$: $r_{i}=b-Kx_{i-1}$, where $x_{i-1}$ is our guess for iteration $i-1$. (Full precision)
  2. 2.
    Solve the linear system under a regularization parameter $(K+\delta I)\Delta_{i}=r_{i}$. (Lower precision)
  3. 3.
    Take $x_{i}=x_{i-1}+\Delta_{i}$. (Full precision)
In all our matrix solves, given matrix $K$ and right-hand side $b$, we factorize the lower precision copy of $K$, and then solve the linear system in this lower precision. A regularization parameter $\delta$ is added to $K$ to increase the numerical stability of matrix factorization. It is the sum of a static regularization and a dynamic regularization. To offset the regularization effect and rounding errors from the lower precision in backsolves, we solve in full precision for the other steps, i.e., we compute the residual $r_{i}$ and save the update $\{x_{i}\}$ in full precision. We loop until either (1) we reach a pre-specified number of max steps for iterative refinement, or (2) the $\ell_{\infty}$ norm of $b-Kx_{i}$ reaches a certain threshold, specifically $\left\|b-Kx_{i}\right\|_{\infty}\leq t_{\text{abs}}+t_{\text{rel}}\left\|b\right\|_{\infty}$. For the mixed precision, we set the static regularization to the square root of machine precision, i.e., $\sqrt{\epsilon}$, and take a maximum of $10$ iterative refinement steps. For the full-precision solve, $t_{\text{abs}}$ (absolute tolerance) and $t_{\text{rel}}$ (relative tolerance) are both $10^{-12}$ by default.
Although the mixed precision for the iterative refinement can improve the numerical stability for solving linear system in lower precision, it is to be expected that the mixed precision may take longer time to converge or fail in cases where the matrix $K$ is extremely ill-conditioned. Caution is required when using mixed precision for numerically hard problems, e.g., conic programs with exponential cones.
4 Numerical experiments
We have benchmarked our Julia GPU implementation of the Clarabel solver [^1] https://github.com/cvxgrp/CuClarabel against the state-of-the-art commercial interior-point solvers Mosek [12] and Gurobi [53]. We also include our Rust CPU implementation of Clarabel solver with the 3rd-party multithreaded supernodal LDL factorization method in the faer-rs package [54]. We have included benchmark results for several classes of problems including quadratic programming (QP), second-order cone programming (SOCP) and exponential cone programming. All benchmarks are performed using the default settings for each solver, with pre-solve disabled where applicable to ensure equivalent problem-solving conditions. No additional iteration limits are imposed beyond each solver’s internal defaults. All experiments were carried out on a workstation with Intel(R) Xeon(R) w9-3475X CPU @ 4.8 GHz with 256 GB RAM and NVIDIA GeForce RTX 4090 24GB GPU. All benchmarks tests are scripted in Julia and access solver interfaces via JuMP [82]. We use Rust compiler version 1.76.0 and Julia version 1.10.2.
4.1 Benchmarking metrics
We choose the same benchmarking tests as used in the Clarabel solver [15], and compare our results using metrics that are commonly used when comparing solve time across different solvers. For a set of $N$ test problems, we define the shifted geometric mean $g_{s}$ as
$g_{s}=\left[\prod_{p=1}^{N}(t_{p,s}+k)\right]^{\frac{1}{N}}-k,$
where $t_{p,s}$ is the time in seconds for solver $s$ to solve problem $p$, and $k=1$ is the shift. The normalized shifted geometric mean is then defined as
$r_{s}=\frac{g_{s}}{\min_{s^{\prime}}g_{s^{\prime}}}.$
Note that the solver with the lowest shifted geometric mean solve time has a normalized score of 1. We assign a solve time $t_{p,s}$ equal to the maximum allowable solve time if solver $s$ fails to solve the problem $p$.
The relative performance ratio for a solver $s$ and a problem $p$ is defined as
$u_{p,s}=\frac{t_{p,s}}{\min_{s^{\prime}}t_{p,s^{\prime}}}.$
The relative performance profile denotes the fraction of problems solved by solver $s$ within a factor $\tau$ of the solve time of the best solver, which is defined as $f^{r}_{s}:\textbf{R}_{+}\mapsto[0,1]$
$f^{r}_{s}(\tau)=\frac{1}{N}\sum_{p}\mathcal{I}_{\leq\tau}(u_{p,s}),$
where $\mathcal{I}_{\leq\tau}(u)=1$ if $\tau\leq u$ and $\mathcal{I}_{\leq\tau}(u)=0$ otherwise. We also compute the absolute performance profile $f^{a}_{s}:\textbf{R}_{+}\mapsto[0,1]$, which denotes the fraction of problems solved by solver $s$ within $\tau$ seconds and is defined as
$f^{a}_{s}(\tau)=\frac{1}{N}\sum_{p}\mathcal{I}_{\leq\tau}(t_{p,s}).$
We showcase plots of performance profiles and tables of shifted geometric means for different classes of problems supported in the Clarabel solver. We included further detailed numerical results for all solvers in §D, including the iteration counts and the total time.
4.2 Quadratic programming
We first present benchmark results for QPs. Note that in this setting, the set $\mathcal{K}$ in ($\mathcal{P}$) is restricted to the composition of zero cones, i.e., linear equality constraints, and nonnegative cones, i.e., linear inequality constraints. We consider two classes of problems, the portfolio optimization problem and the Huber fitting problem.
Portfolio optimization, a problem arising in quantitative finance, aims to allocate assets in a manner that maximizes expected return while keeping risk under control. We can formulate it as
$\begin{array}[]{ll}\text{maximize}&\mu^{T}x-\gamma x^{T}\Sigma x\\ \text{subject to}&\mathbf{1}^{T}x=1,\\ &x\geq 0,\end{array}$
where $x\in\textbf{R}^{n}$ (the variable) represents the ratio of allocated assets, $\mu\in\textbf{R}^{n}$ is the vector of expected returns, $\gamma>0$ is the risk-aversion parameter, and $\Sigma\in\mathbf{S}_{+}^{n}$ the risk covariance matrix which is of the form $\Sigma=FF^{\top}+D$ with $F\in\textbf{R}^{p\times n}$ and $D\in\textbf{R}^{p\times p}$ diagonal. We set the rank $p$ to the integer closest to $0.1n$, and vary $n$ from $5000$ to $25000$.
Huber fitting is a version of robust least squares. For a given matrix $A\in\textbf{R}^{m\times n}$ and vector $b\in\textbf{R}^{m}$, we replace the least squares loss function with the Huber loss. The Huber loss makes the penalty incurred by larger points linear instead of quadratic, thus outliers have a smaller effect on the resulting estimator. Precisely, the problem is stated as:
$\text{minimize}\ \sum_{i=1}^{m}\phi_{h}\left(a_{i}^{T}x-b_{i}\right)$
where $a_{i}^{\top}$ is the $i$-th row of $A$ and the Huber loss $\phi_{h}:\textbf{R}\rightarrow\textbf{R}$ is defined as
$\phi_{h}(t)=\begin{cases}t^{2}&|t|\leq T\\ T(2|t|-T)&\text{ otherwise }\end{cases}.$
We set $m$ to the nearest integer of $1.5n$ and vary value of $n$ from $5000$ to $25000$.
Figure 2: Performance profiles for the large QPs problem set
(a) Relative performance profile
(b) Absolute performance profile
ClarabelGPU ClarabelRs Gurobi Mosek
Shifted GM Full Acc. 1.0 11.29 2.13 4.18
Low Acc. 1.0 11.29 2.13 4.18
Failure Rate (%) Full Acc. 0.0 20.0 0.0 0.0
Low Acc. 0.0 10.0 0.0 0.0
(c) Benchmark timings as shifted geometric mean and failure rates
Results for large QP tests are shown in Figure 2. We benchmark ten different examples from two classes above and set the time limit to $1$h. We compare our GPU implementation ClarabelGPU with ClarabelRs, the Rust version 0.9.0 version of Clarabel equipped with the faer multithreaded linear system solver; and two commercial solvers, Gurobi and Mosek. ClarabelGPU is the fastest solver on these problems, and it has the lowest per-iteration time for almost all examples. Since most of time of an interior point solver is spent on factorizing and solving a linear system in QPs, we can say that ClarabelGPU benefits from the use of the CUDSS linear system solver and it is more than 2 times faster than Gurobi, about 4 times faster than Mosek and 10x times faster than the existing Rust implementation with the multithreaded faer linear system solver.
4.3 Second-order cone programming
We next consider the second-order cone relaxations of optimal power flow problems [55] from the IEEE PLS PGLib-OPF benchmark library [2], using the PowerModels.jl package [56] for modeling convenience. Note that only second-order cones of dimensionality $3$ or $4$ are used in these second-order cone relaxations, which satisfies our assumption in §3.2: that the dimensionality of each cone is very small.
We compare our GPU implementation with the CPU-based Clarabel solver and Mosek solver. We also include results of the Mosek solver with pre-solve for comparison, which is denoted as Mosek* in the plots. The maximum termination time is again set to $1$h. We benchmark the second-order cone relaxations of 120 problems from the PGLib-OPF library, where the number of second-order cones exceed $2000$.
Results for these problems are shown in Figure 3. Both ClarabelGPU and ClarabelRs are faster and more numerically stable than Mosek even with the presolve step. Moreover, the GPU implementation can solve 118 out of 120 examples within an hour, while the Rust version of Clarabel can solve 104 out of 120 examples within the same time limit. In contrast, Mosek with presolve fails on about 40% of the optimal flow problems, and the one without presolve fails on over 90% of the problems.
Overall, the GPU solver is several times faster than Rust-based CPU solver. However, note that ClarabelGPU fails on two examples that ClarabelRs successfully solves. This highlights the different numerical performance of the linear system solvers between CUDSS and faer.
Figure 3: Performance profiles for the large OPF SOCPs problem set
(a) Relative performance profile
(b) Absolute performance profile
ClarabelGPU Mosek* ClarabelRs Mosek
Shifted GM Full Acc. 1.0 6.63 3.62 34.33
Low Acc. 1.0 7.02 2.99 36.36
Failure Rate (%) Full Acc. 1.7 39.0 13.6 90.7
Low Acc. 0.0 39.0 3.4 90.7
(c) Benchmark timings as shifted geometric mean and failure rates
4.4 Exponential cone programming
We benchmark two classes of exponential cone programming problems with varying dimensionality, the logistic regression and the entropy maximization problems [13].
Logistic regression is a classical machine learning model for binary classification. Given a feature matrix $A\in\textbf{R}^{m\times n}$ and a vector of labels $b\in\textbf{R}^{m}$ such that $b_{i}\in\{0,1\}$, we formulate the sparse logistic regression problem as:
$\text{minimize}\ \ \sum_{i=1}^{m}\left(\log\left(1+\exp\left(x^{T}a_{i}\right)\right)-b_{i}x^{T}a_{i}\right)+\lambda\|x\|_{1},$
where $x$ is our variable and $a_{i}^{T}$ is the $i$-th row of $A$. Note that this is the sparse logistic regression problem due to the $\ell_{1}$-norm penalty on $x$. Indeed, when $\lambda=0$, we recover the classical logistic regression problem.
Each entry of $A$ is generated via $A_{ij}\sim\mathcal{N}(0,1)$. The label $b$ is generated following the same method in [57, 13]. We set $m=5n$ and vary value of $n$ from $2000$ to $10000$.
The entropy maximization problem aims to maximize entropy over a probability distribution given a set of $m$ linear inequality constraints, which can be interpreted as bounds on the expectations of arbitrary functions. The problem is formulated as
$\begin{array}[]{ll}\text{maximize}&-\sum_{i=1}^{n}x_{i}\log x_{i}\\ \text{subject to}&1^{T}x=1,\\ &Ax\leq b,\end{array}$
Each element of $A$ is generated from the distribution $A_{ij}\sim\mathcal{N}(0,n)$. Then, we set $b=Av/1^{\top}v$ where $v\in\textbf{R}^{n}$ is generated randomly from $v_{i}\sim U[0,1]$, which ensures the problem is always feasible. We set $m$ to the nearest integer of $0.5n$ and vary value of $n$ from $2000$ to $10000$.
The benchmark results for these two problems with varying dimensionality are shown in Figure 4. From §D, even though the GPU acceleration of Clarabel is offset by nearly doubled number of iterations compared to Mosek, we can still achieve more than 2 times of acceleration for the overall time. That is to say we can possibly achieve more acceleration in ClarabelGPU if we can improve the numerical stability to the same level of Mosek on exponential cone programs. Compared to ClarabelRs, we find ClarabelGPU can benefit from GPU computation up to 20x times faster on logistic regression problems.
Figure 4: Performance profiles for the exponential cone programming problem set
(a) Relative performance profile
(b) Absolute performance profile
ClarabelGPU Mosek* ClarabelRs Mosek
Shifted GM Full Acc. 1.0 1.47 8.42 1.58
Low Acc. 1.0 1.47 8.42 1.58
Failure Rate (%) Full Acc. 0.0 0.0 40.0 0.0
Low Acc. 0.0 0.0 20.0 0.0
(c) Benchmark timings as shifted geometric mean and failure rates
4.5 Mixed precision
In the implementation of the mixed precision setting, as described we only set the data type of the CUDSS linear system solver to Float32. Since the use of mixed precision accelerates the numerical factorization rather than the symbolic factorization within a factorization method, we only record the computational time for an interior point method without the setup time.
We test the mixed precision setting on QPs, including portfolio optimization and Huber fitting problems from §4.2. We compare it with the standard GPU implementation of Float64 data type. The results in Table LABEL:table:mixed_large_qp show that the use of mixed precision strategy can reduce the solve time up to a factor of $2$. This will be beneficial to problems like the parametric programming, where the setup is needed only once and the symbolic factorization structure can be reused repeatedly later on [58]. However, we also notice that the mixed precision may not be sufficient for conic optimization problems where the Hessian block $H$ in (7) is ill-conditioned, especially for the exponential cone that the exponential component causes huge variation in the magnitude of different entries.
Table 1: Solve times and iteration counts for the mixed precision QP test
iterations total time (s)
Problem Full Mixed Full Mixed
portfolio_optimization_n_10000 19 19 0.729 0.631
portfolio_optimization_n_25000 19 21 7.08 4
portfolio_optimization_n_20000 19 23 3.91 2.54
huber_fitting_n_5000 9 9 2.61 1.32
huber_fitting_n_25000 10 10 143 58.3
huber_fitting_n_10000 10 10 12.6 5.76
portfolio_optimization_n_5000 17 18 0.196 0.259
huber_fitting_n_15000 10 10 34.9 15.1
portfolio_optimization_n_15000 19 19 1.88 1.18
huber_fitting_n_20000 9 9 69.8 29.1
5 Conclusion
We have developed a GPU interior point solver for conic optimization. [^2] https://github.com/cvxgrp/CuClarabel We propose a mixed parallel computing strategy to process linear constraints with second-order cone, exponential cone and power cone constraints. The GPU solver shows several times of acceleration compared to state-of-the-art CPU conic solvers on many problems to high precision, including QPs, SOCPs and exponential programmings.
We also note that some initialization steps, such as population of the KKT matrix and matrix equilibration, take non-negligible time for large problems. We will also move the setup steps onto a GPU in our future development. The proposed mixed parallel computing strategy for GPU implementation is also applicable for conic solvers based on first-order operator-splitting methods.
Acknowledgments
The work of Yuwen Chen was funded by EPSRC Impact Acceleration Account Fund. Parth Nobel was supported in part by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE-1656518. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. The work of Stephen Boyd was supported by ACCESS (AI Chip Center for Emerging Smart Systems), sponsored by InnoHK funding, Hong Kong SAR.
Appendix A Supported cones in Clarabel and CuClarabel
We support the following atomic cones in our GPU solver:
  • The zero cone, defined as
    $\{0\}^{n}=\left\{x\in\textbf{R}^{n}\ \middle|\ x_{i}=0,~{}i=1,\ldots,n\right\}.$
    The dual cone of the zero cone is ($\left(\{0\}^{n}\right)^{*}=\textbf{R}^{n}$.
  • The nonnegative cone, defined as
    $\textbf{R}_{+}^{n}=\left\{x\in\textbf{R}^{n}\ \middle|\ x_{i}\geq 0,~{}i=1,\ldots,n\right\}.$
    The nonnegative cone is a self-dual convex cone, i.e., $\left(\textbf{R}_{+}^{n}\right)^{*}=\textbf{R}_{+}^{n}$.
  • The second-order cone $\mathcal{K}_{\text{soc }}^{n}$ (also called the quadratic or Lorentz cone), defined as
    $\mathcal{K}_{\mathrm{soc}}^{n}=\left\{(t,x)\ \middle|\ x\in\textbf{R}^{n-1},~{}t\in\textbf{R}_{+},~{}\|x\|_{2}\leq t\right\}.$
    The second-order cone is self-dual, i.e., $\mathcal{K}_{\mathrm{soc}}=\mathcal{K}_{\mathrm{soc}}^{*}$.
  • The exponential cone, a $3$-dimensional cone defined as
    $\mathcal{K}_{\exp}=\left\{(x,y,z)\ \middle|\ y>0,~{}y\exp\left(\frac{x}{y}\right)\leq z\right\}\cup\{(x,0,z)\mid x\leq 0,~{}z\geq 0\},$
    with its dual cone given by
    $\mathcal{K}_{\exp}^{*}=\left\{(u,v,w)\ \middle|\ u<0,~{}-u\exp\left(\frac{v}{u}\right)\leq\exp(1)w\right\}\cup\{(0,v,w)\mid v\geq 0,~{}w\geq 0\}.$
  • The $3$-dimensional power cone with exponent $\alpha\in(0,1)$, defined as
    $\mathcal{K}_{\text{pow},\alpha}=\left\{(x,y,z)\ \middle|\ x^{\alpha}y^{1-\alpha}\geq|z|,~{}x\geq 0,~{}y\geq 0\right\},$
    with its dual cone given by
    $\mathcal{K}_{\text{pow},\alpha}^{*}=\left\{(u,v,w)\ \middle|\ \left(\frac{u}{\alpha}\right)^{\alpha}\left(\frac{v}{1-\alpha}\right)^{1-\alpha}\geq|w|,~{}u\geq 0,~{}v\geq 0\right\}.$
Appendix B Barrier functions for different cones
For cones introduced in §A, the corresponding logarithmically-homogeneous self-concordant barrier functions with degree $\nu$ are given by:
  1. 1.
    Nonnegative cone $\textbf{R}_{+}^{n}$ of degree $n$:
    $f(z)=-\sum_{i\in\llbracket n\rrbracket}\log(z_{i}),\ z\in\textbf{R}_{+}^{n}.$
  2. 2.
    Second order cone $\mathcal{K}_{q}^{n}$ of degree $1$:
    $f(z)=-\frac{1}{2}\log\left(z_{1}^{2}-\sum_{i=2}^{n}z_{i}^{2}\right),\ z\in\mathcal{K}_{q}^{n}.$
  3. 3.
    Dual exponential cone $\mathcal{K}_{\mathrm{exp}}^{*}$ of degree $3$:
    $f(z)=-\log\left(z_{2}-z_{1}-z_{1}\log\left(\frac{z_{3}}{-z_{1}}\right)\right)-\log(-z_{1})-\log(z_{3}),\ z\in\mathcal{K}_{\mathrm{exp}}^{*}.$
  4. 4.
    Dual power cone $\mathcal{K}_{\mathrm{pow}}^{*}$ of degree $3$:
    $f(z)=-\log\left(\left(\frac{z_{1}}{\alpha}\right)^{2\alpha}\left(\frac{z_{2}}{1-\alpha}\right)^{2(1-\alpha)}-z_{3}^{2}\right)-(1-\alpha)\log(z_{1})-\alpha\log(z_{2}),\ z\in\mathcal{K}_{\mathrm{pow}}^{*}.$
Appendix C Scaling matrices
The scaling matrix $H$ in (6) is the Jacobian from the linearization of the central path (5). We set $H=0$ for the zero cone and choose the NT scaling [51] for nonnegative and second-order cones, which are both symmetric cones. The NT scaling method exploits the self-scaled property of symmetric cone $\mathcal{K}$ where exists a unique scaling point $w\in\mathcal{K}$ satisfying
$H(w)s=z.$
The matrix $H(w)$ can be factorized as $H^{-1}(w)=W^{\top}W$, and we set $H=H^{-1}(w)$ in (6). The factors $w,W$ are then computed following [47].
For exponential and power cones that are not symmetric, the central path is defined by the set of point satisfying
$Hz=s,\quad H\nabla f^{*}(s)=\nabla f(z),$
where $f^{*}$ is the conjugate function of $f$, and the symmetric scaling from [59, 50] is implemented. We define shadow iterates as
$\tilde{z}=-\nabla f(s),\quad\tilde{s}=-\nabla f^{*}(z),$
with $\tilde{\mu}=\langle\tilde{s},\tilde{z}\rangle/\nu$. A scaling matrix $H$ is chosen to be the rank-4 Broyden-Fletcher-Goldfarb-Shanno (BFGS) update as in a quasi-Newton method,
$H=H_{\mathrm{BFGS}}=Z(Z^{\top}S)^{-1}Z^{\top}+H_{a}-H_{a}S(S^{\top}H_{a}S)^{-1}S^{\top}H_{a},$
where $Z=[z,\tilde{z}],S=[s,\tilde{s}]$ and $H_{a}=\mu\nabla^{2}f(z)$ is an estimate of the Hessian as in [50].
Algorithm 1 shows how to update the scaling matrix $H$ at iteration $k+1$:
Algorithm 1 Update the scaling matrix $H$
1:  Input: the scaling matrix $H^{k}$ at iteration $k$
2:
3:  Update $H_{1}^{k+1}$   // Zero cone
4:  Update $H_{2}^{k+1}$   // Nonnegative cone
5:
6:  @sync for $t=3$ to $i$   // Second-order cones
7:  $H_{t}^{k+1}=$soc_update_H($H_{t}^{k}$)
8:  end
9:
10:  @sync for $t=i+1$ to $j$   // Exponential cones
11:  $H_{t}^{k+1}=$exp_update_H($H_{t}^{k}$)
12:  end
13:
14:  @sync for $t=j+1$ to $p$   // Power cones
15:  $H_{t}^{k+1}=$pow_update_H($H_{t}^{k}$)
16:  end
17:
18:  return  $H^{k+1}$   // Output the new scaling matrix $H^{k+1}$
Appendix D Detailed benchmark results
Table 2: Solve times and iteration counts for the large LPs problem set
iterations total time (s)
Problem vars. cons. nnz(A) nnz(P) ClarabelGPU Mosek ClarabelGPU Mosek
qap15 28605 22275 117225 0 22 16 4.72 3.6
brazil3 62488 23968 181026 0 20 15 0.66 1.53
ex10 104968 17680 1197360 0 13 16 5.75 14.7
neos_5052403_cygnet 104004 32868 4964040 0 23 12 7.66 4.86
supportcase10 195224 14770 584622 0 35 20 9.6 5.65
square41 164591 62234 13690857 0 28 12 8.39 28.5
chromaticindex1024_7 215039 73728 417780 0 9 10 1.8 6.4
neos_3025225 231263 69846 9497642 0 28 19 50.4 19.9
rmine15 443271 42438 964608 0 55 29 41.2 14
physiciansched3_3 417922 79555 1214174 0 452 284 251 174
neos_5251015 486996 136971 1955853 0 - 9 - 22.6
graph40_40 566100 102600 1466100 0 11 - 3.36 -
datt256_lp 535365 262144 2028020 0 10 5 7.52 223
s250r10 557246 273142 1864891 0 103 32 29.3 6.56
s100 743567 364417 2506751 0 92 33 96.1 9.42
savsched1 877295 328575 2351813 0 28 16 10.9 85.8
woodlands09 958865 382147 3410269 0 - 8 - 344
scpm1 1005000 500000 7250000 0 27 24 24.4 13.2
bharat 2030687 590519 4296196 0 67 - 67.3 -
tpl_tub_ws1617 2649817 747601 6215769 0 141 63 110 133
fhnw_binschedule1 3056177 1141653 10894631 0 49 - 232 -
supportcase19 2868909 1429098 7145290 0 64 17 56 12.5
s82 3469140 1690631 10403870 0 - 71 - 235
Table 3: Solve times and iteration counts for the QPs problem set
iterations total time (s)
Problem vars. cons. nnz(A) nnz(P) ClarabelGPU ClarabelRs Gurobi Mosek ClarabelGPU ClarabelRs Gurobi Mosek
portfolio_optimization_n_5000 5501 5500 1259971 5500 15 15 17 9 0.442 7.9 0.9 2.65
portfolio_optimization_n_10000 11001 11000 5022139 11000 18 18 19 7 1.67 57.1 4.53 10.7
portfolio_optimization_n_15000 16501 16500 11282198 16500 18 18 18 8 4.16 165 10.4 28.7
portfolio_optimization_n_20000 22001 22000 20040608 22000 17 17 18 8 8.06 372 22.3 63.1
huber_fitting_n_5000 22500 27500 4725756 7500 8 8 13 15 4.24 68.7 6.53 10.3
portfolio_optimization_n_25000 27501 27500 31300056 27500 18 18 20 9 14.7 600 44.2 124
huber_fitting_n_10000 45000 55000 18826645 15000 8 8 11 17 18.7 636 39.7 77
huber_fitting_n_15000 67500 82500 42304462 22500 8 8 11 18 48.6 2.18e+03 110 383
huber_fitting_n_20000 90000 110000 75138584 30000 8 - 14 20 99.2 - 269 1.21e+03
huber_fitting_n_25000 112500 137500 117388752 37500 8 - 11 19 180 - 400 2.31e+03
Table 4: Solve times and iteration counts for the large OPF LP problem set
iterations total time (s)
Problem vars. cons. nnz(A) nnz(P) ClarabelGPU ClarabelRs Mosek ClarabelGPU ClarabelRs Mosek
case10192_epigrids 92051 27914 180020 697 17 17 35 0.742 0.533 0.764
case10192_epigrids__api 92051 27914 180020 697 17 17 28 0.706 0.526 0.627
case10480_goc 99926 29816 196673 276 20 20 27 0.82 0.712 0.701
case10480_goc__api 99926 29816 196673 276 18 18 31 0.806 0.666 0.75
case13659_pegase 120495 38218 230046 0 122 107 81 2.21 3.29 2.69
case13659_pegase__api 120495 38218 230046 0 15 15 59 0.876 0.648 1.82
case20758_epigrids 182928 56275 355508 1881 18 18 32 1.43 1.43 1.28
case20758_epigrids__api 182928 56275 355508 1881 18 18 31 1.4 1.41 1.25
case19402_goc 184959 55077 364846 249 24 24 41 1.65 1.86 1.65
case19402_goc__api 184959 55077 364846 249 25 25 38 1.72 2.03 1.59
case19402_goc__sad 184959 55077 364846 249 22 - 55 1.61 - 2.09
case24464_goc 210481 63871 408258 348 22 - 26 1.82 - 1.52
case24464_goc__api 210481 63871 408258 348 21 21 25 1.76 2.09 1.41
case24464_goc__sad 210481 63871 408258 348 24 - 65 1.82 - 2.74
case30000_goc 213698 68919 399262 372 38 38 47 1.99 2.83 1.45
case30000_goc__api 213698 68919 399262 372 30 30 39 1.91 2.54 1.34
case30000_goc__sad 213698 68919 399262 372 25 25 83 1.65 1.89 2.24
case78484_epigrids 685984 211266 1334253 0 29 29 117 8.31 16.8 68.9
case78484_epigrids__api 685984 211266 1334253 0 26 26 106 7.89 15.2 93.8
case78484_epigrids__sad 685984 211266 1334253 0 30 30 120 8.37 17.6 68.7
Table 5: Solve times and iteration counts for the large OPF SOCP problem set
iterations total time (s)
Problem vars. cons. nnz(A) nnz(P) ClarabelGPU Mosek* ClarabelRs Mosek ClarabelGPU Mosek* ClarabelRs Mosek
case2383wp_k 97600 20393 155950 0 109 - 107 - 1.81 - 3.63 -
case2383wp_k__api 97600 20393 155950 0 21 38 23 33 0.928 1.21 1.26 1.89
case2383wp_k__sad 97600 20393 155950 0 99 - 99 - 1.78 - 3.48 -
case2312_goc 98556 20518 159516 0 47 151 48 - 1.35 3.98 2.05 -
case2312_goc__api 98556 20518 159516 0 65 250 65 - 1.41 6.23 2.19 -
case2312_goc__sad 98556 20518 159516 0 59 191 59 - 1.32 4.98 1.98 -
case2737sop_k 109822 22777 176602 0 70 - 75 - 1.6 - 3.07 -
case2737sop_k__api 109822 22777 176602 0 71 - 73 - 1.63 - 2.99 -
case2737sop_k__sad 109822 22777 176602 0 72 - 73 - 1.62 - 2.92 -
case2736sp_k 110022 22878 176901 0 71 - 69 - 1.63 - 2.88 -
case2736sp_k__api 110022 22878 176901 0 57 83 57 - 1.41 2.63 2.65 -
case2736sp_k__sad 110022 22878 176901 0 71 - 72 - 1.53 - 3.67 -
case2746wp_k 111106 23320 178556 0 74 - 73 - 1.78 - 3.28 -
case2746wp_k__api 111106 23320 178556 0 31 33 32 35 1.18 1.22 1.78 2.36
case2746wp_k__sad 111106 23320 178556 0 74 133 74 - 1.65 4.28 3.19 -
case2746wop_k 111822 23434 179829 0 64 - 64 - 1.57 - 2.8 -
case2746wop_k__api 111822 23434 179829 0 26 33 - 36 1.11 1.18 - 2.21
case2746wop_k__sad 111822 23434 179829 0 64 - 63 - 1.43 - 2.6 -
case3012wp_k 120676 25202 193907 0 94 - 91 - 1.96 - 3.94 -
case3012wp_k__api 120676 25202 193907 0 29 53 - 48 1.2 1.98 - 3.1
case3012wp_k__sad 120676 25202 193907 0 89 - 91 - 1.94 - 3.92 -
case2848_rte 122708 25858 197228 0 80 177 80 - 2.15 5.3 4.75 -
case2848_rte__api 122708 25858 197228 0 75 - 75 - 2.28 - 5.25 -
case2848_rte__sad 122708 25858 197228 0 70 - 70 - 2.39 - 4.36 -
case2868_rte 123912 26164 198869 0 84 225 84 - 2.35 6.49 5.8 -
case2868_rte__api 123912 26164 198869 0 73 258 75 - 2.11 7.53 4.96 -
case2868_rte__sad 123912 26164 198869 0 79 288 79 - 2.24 8.31 5.24 -
case3120sp_k 124354 25856 199827 0 79 - 75 - 1.77 - 3.4 -
case3120sp_k__api 124354 25856 199827 0 87 - 92 - 2.02 - 4.14 -
case3120sp_k__sad 124354 25856 199827 0 99 - 100 - 2.03 - 4.4 -
case2853_sdet 128886 27445 206896 0 102 269 103 - 2.1 8.51 4.34 -
case2853_sdet__api 128886 27445 206896 0 122 309 121 - 2.43 9.57 5.12 -
case2853_sdet__sad 128886 27445 206896 0 101 276 101 - 2.1 8.64 4.33 -
case3022_goc 134780 28060 217434 0 41 174 - 139 1.52 6.16 - 9.67
case3022_goc__api 134780 28060 217434 0 58 202 - 177 1.84 6.88 - 12.9
case3022_goc__sad 134780 28060 217434 0 42 177 - 142 1.51 5.82 - 9.91
case3375wp_k 139126 29112 223999 0 101 - 89 - 2.14 - 4.04 -
case3375wp_k__api 139126 29112 223999 0 146 - 151 - 2.81 - 6.84 -
case3375wp_k__sad 139126 29112 223999 0 103 - 92 - 2.17 - 4.25 -
case2869_pegase 143608 30153 236217 0 53 235 53 - 2.03 8.56 4.44 -
case2869_pegase__api 143608 30153 236217 0 50 264 50 - 1.88 9.44 4.4 -
case2869_pegase__sad 143608 30153 236217 0 48 278 48 - 1.86 9.99 4.08 -
case2742_goc 144110 29856 236467 0 110 119 109 - 3.89 4.9 9.42 -
case2742_goc__api 144110 29856 236467 0 104 127 102 129 3.46 5.25 8.19 10.4
case2742_goc__sad 144110 29856 236467 0 107 119 106 - 3.83 4.94 9.09 -
case4661_sdet 198498 41599 320728 0 140 - 137 - 4.23 - 11.2 -
case4661_sdet__api 198498 41599 320728 0 125 - 125 - 3.94 - 10.2 -
case4661_sdet__sad 198498 41599 320728 0 139 - 134 - 4.37 - 10.9 -
case3970_goc 205819 42789 337328 0 140 154 143 - 4.94 9.11 15.6 -
case3970_goc__api 205819 42789 337328 0 109 136 115 - 4.85 8.04 15.1 -
case3970_goc__sad 205819 42789 337328 0 142 169 144 - 5.07 10.2 16.2 -
case4020_goc 216455 44877 355706 0 181 194 186 - 6.86 12.1 25.1 -
case4020_goc__api 216455 44877 355706 0 101 144 101 - 4.77 9.28 15.4 -
case4020_goc__sad 216455 44877 355706 0 183 178 185 - 6.71 13.2 24.4 -
case4917_goc 218213 45522 352833 0 52 164 - 147 2.99 8.63 - 16.7
case4917_goc__api 218213 45522 352833 0 62 260 60 - 3.27 13.4 7.7 -
case4917_goc__sad 218213 45522 352833 0 47 184 - 166 2.91 10.1 - 18.8
case4601_goc 225588 46885 368445 0 179 192 180 - 5.98 11.5 21.8 -
case4601_goc__api 225588 46885 368445 0 103 147 109 162 4.66 8.99 15.6 20.5
case4601_goc__sad 225588 46885 368445 0 180 191 177 - 5.94 12.2 20 -
case4837_goc 240160 49855 392884 0 139 152 137 - 7.18 10.6 24.3 -
case4837_goc__api 240160 49855 392884 0 122 177 124 - 5.95 12.1 19 -
case4837_goc__sad 240160 49855 392884 0 136 161 130 - 7.05 11 23 -
case4619_goc 254753 52616 419210 0 145 153 144 - 7.98 11 26.8 -
case4619_goc__api 254753 52616 419210 0 130 161 130 - 6.85 11.9 21.4 -
case4619_goc__sad 254753 52616 419210 0 151 150 151 - 8.17 11.2 27 -
case5658_epigrids 282180 58620 461724 0 150 181 143 - 9.01 13.8 28.4 -
case5658_epigrids__api 282180 58620 461724 0 164 202 159 - 10.8 15 34.9 -
case5658_epigrids__sad 282180 58620 461724 0 156 221 158 - 8.82 16.6 29.1 -
case6468_rte 286248 59396 463599 0 104 250 101 - 5.5 16.1 17.7 -
case6468_rte__api 286248 59396 463599 0 88 292 88 - 5.43 18.4 17.2 -
case6468_rte__sad 286248 59396 463599 0 93 254 99 - 5.2 16.5 17.5 -
case6470_rte 287806 60144 465757 0 122 319 109 - 6.14 20.3 21.9 -
case6470_rte__api 287806 60144 465757 0 86 330 85 - 5.22 21.4 17.2 -
case6470_rte__sad 287806 60144 465757 0 102 350 99 - 5.49 22.2 17.5 -
case6495_rte 288050 60099 465939 0 112 329 115 - 6.07 21.3 20.1 -
case6495_rte__api 288050 60099 465939 0 102 349 95 - 5.61 22.3 16.6 -
case6495_rte__sad 288050 60099 465939 0 112 344 108 - 5.64 22.1 18.7 -
case6515_rte 288710 60239 466922 0 97 299 95 - 6.3 19.1 17.7 -
case6515_rte__api 288710 60239 466922 0 81 332 81 - 5.29 21.7 15.8 -
case6515_rte__sad 288710 60239 466922 0 86 327 84 - 5.24 21.4 15.6 -
case7336_epigrids 358450 74618 585766 0 156 196 154 - 11.6 18 47.8 -
case7336_epigrids__api 358450 74618 585766 0 150 214 152 - 10.5 19.9 43.6 -
case7336_epigrids__sad 358450 74618 585766 0 155 186 156 - 11.5 17.3 50.2 -
case10000_goc 440997 92799 712177 0 - 86 - - - 10.3 - -
case10000_goc__api 440997 92799 712177 0 62 206 - - 6.91 21.1 - -
case10000_goc__sad 440997 92799 712177 0 90 152 - - 7.44 16.4 - -
case8387_pegase 459046 96351 750588 0 139 - 138 - 9.19 - 30.3 -
case8387_pegase__api 459046 96351 750588 0 140 - 140 - 10 - 34.3 -
case8387_pegase__sad 459046 96351 750588 0 134 - 134 - 9.01 - 29.8 -
case9591_goc 495073 102120 811889 0 226 - 229 - 21 - 85.7 -
case9591_goc__api 495073 102120 811889 0 149 159 149 - 15.1 25.5 58.6 -
case9591_goc__sad 495073 102120 811889 0 228 200 228 - 21.3 31.5 86.8 -
case9241_pegase 502110 104741 827419 0 70 - 66 - 6.83 - 20.3 -
case9241_pegase__api 502110 104741 827419 0 67 - 70 - 6.61 - 20.6 -
case9241_pegase__sad 502110 104741 827419 0 70 389 68 - 6.73 58.4 20.5 -
case10192_epigrids 529695 109758 866501 0 96 113 96 - 10.6 17 36 -
case10192_epigrids__api 529695 109758 866501 0 91 113 91 - 10.4 17.4 35.3 -
case10192_epigrids__sad 529695 109758 866501 0 109 - 107 - 11.9 - 41.5 -
case10480_goc 573754 118760 943278 0 164 - 162 - 18.2 - 76.2 -
case10480_goc__api 573754 118760 943278 0 136 217 - - 16.2 40.5 - -
case10480_goc__sad 573754 118760 943278 0 164 242 164 - 18.3 43.6 77.4 -
case13659_pegase 662910 140961 1081042 0 83 - 82 - 11.8 - 42.9 -
case13659_pegase__api 662910 140961 1081042 0 78 - 78 - 11.4 - 39.9 -
case13659_pegase__sad 662910 140961 1081042 0 78 329 - - 11 60 - -
case20758_epigrids 1048059 218151 1709288 0 122 - 123 - 27.5 - 99.1 -
case20758_epigrids__api 1048059 218151 1709288 0 154 - 155 - 39.9 - 150 -
case20758_epigrids__sad 1048059 218151 1709288 0 137 - 135 - 30.3 - 111 -
case19402_goc 1064421 219911 1753874 0 256 - 251 - 45.9 - 221 -
case19402_goc__api 1064421 219911 1753874 0 212 - 216 - 39.3 - 193 -
case19402_goc__sad 1064421 219911 1753874 0 257 - 258 - 46 - 232 -
case30000_goc__api 1195834 249462 1921390 0 150 - - - 24.6 - - -
case24464_goc 1202964 248644 1983519 0 230 - - - 39.4 - - -
case24464_goc__api 1202964 248644 1983519 0 213 - 214 - 33.4 - 149 -
case24464_goc__sad 1202964 248644 1983519 0 - - 224 - - - 169 -
case78484_epigrids 3904758 811998 6389202 0 401 - - - 260 - - -
case78484_epigrids__api 3904758 811998 6389202 0 400 - 401 - 320 - 1.65e+03 -
case78484_epigrids__sad 3904758 811998 6389202 0 408 - - - 263 - - -
Table 6: Solve times and iteration counts for the exponential cone programming problem set
iterations total time (s)
Problem vars. cons. nnz(A) nnz(P) ClarabelGPU Mosek* ClarabelRs Mosek ClarabelGPU Mosek* ClarabelRs Mosek
entropy_N_2000 11001 4000 258917 0 23 15 21 10 0.548 0.506 3.05 0.565
entropy_N_4000 22001 8000 1019263 0 24 16 20 9 1.87 2.29 14.2 2.22
entropy_N_6000 33001 12000 2278947 0 26 16 22 9 3.56 6.09 45.1 5.85
entropy_N_8000 44001 16000 4037959 0 27 16 22 10 6.2 12.3 107 12.6
entropy_N_10000 55001 20000 6297376 0 27 16 22 10 9.96 22.6 235 40.1
log_reg_n_2000 94002 34001 1320255 0 57 38 121 37 4.13 3.96 107 5.43
log_reg_n_4000 188002 68001 5205338 0 68 47 - 47 17.4 16.4 - 21.6
log_reg_n_6000 282002 102001 11529263 0 72 49 - 51 40.9 67.7 - 59.8
log_reg_n_8000 376002 136001 20589037 0 73 48 - 50 77.5 135 - 105
log_reg_n_10000 470002 170001 31904473 0 84 54 - 58 150 372 - 336